OpenStudy (anonymous):

find the area of the region bounded by the functions f(x)=-x^2 +9x-4, x=-1, x=2, and y=0

OpenStudy (anonymous):

x=-1. are you sure it isn't x=1?

OpenStudy (anonymous):

yes im sure it is x=-1

OpenStudy (anonymous):

i don't think there can be an area if y=0

OpenStudy (anonymous):

that is why was not sure. so the answer is no area?

OpenStudy (anonymous):

maybe, if you were not given that value however, you would integrate with the x values acting as limits.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[\begin{align*}f(x)&=-x^2+9x-4\\\\ &=-\left(x^2-9x\right)-4\\\\ &=-\left(x^2-9x+\frac{81}{4}-\frac{81}{4}\right)-4\\\\ &=-\left(\left(x-\frac{9}{2}\right)^2-\frac{81}{4}\right)-4\\\\ &=-\left(x-\frac{9}{2}\right)^2+\frac{81}{4}-4\\\\ &=-\left(x-\frac{9}{2}\right)^2+\frac{65}{4} \end{align*}\]  There's definitely an area...